# Predicting New Research Directions in Materials Science using Large Language Models and Concept Graphs

### v1

@@ Abstract
Table S11: Concept combinations ranked during the interviews (25 examples per category) (continued)
Figure S8: Overview of the concept categories per interview. A1 and A2 are grouped together. Note that the results from researcher 3 were obtained by a commentary on the report, not by an interview.
Human expert evaluation: Examples
To illustrate the value of the suggested combinations of concepts in more detail, we present below five cherry-picked examples by putting them into context and describing in more detail why these combinations are interesting.
@@
[…]
@@
[…]
@@
Recent research underscores that molecular tensile strain management might be regarded as a more general proxy for interface design. Brabec and Friederich reported in Science 2024 [wu2024inverse] torsional flexibility of HTL molecules, which was represented by the introduction of TPA units, as one of several design criteria for efficient perovskite solar cells. We note that the suggestion "tensile strain + molecular architecture" was generated in December 2023. Properties such as film formation and homogenous film quality benefited from that as well. In another currently ongoing study by Brabec et al. (in submission), we found that embedding plasticizing molecules dramatically increases the mechanical and operational stability of thin film solar cells.
@@
The link between stress-induced phase transformation and hexagonal boron nitride (h-BN) is best explained in the context of enhancing the fracture toughness of ceramics by phase transformations. For instance, in zirconia, a tetragonal-to-monoclinic phase transformation results in a volume expansion in the vicinity of a crack, which locally results in compressive stress both, deflecting and stopping crack growth. This concept is applied to other metastable phases in hard coatings [Wang2014]. A second concept of reaching an enormous tensile strength with simultaneously high fracture toughness, which both typically exclude each other, is using local anisotropy. This was shown in pearlite wires, where multiple micro-cracks parallel to the loading direction were formed and enhanced the fracture toughness substantially [Hohenwarter2016]. There are at least two arguments why thinking of stress-induced phase transformations and h-BN is reasonable. (i) Cubic boron nitride (c-BN), having a Wurtzite crystal structure, is used as a hard coating because of its enormous hardness. In contrast, the hardness of h-BN is significantly lower, rendering it not a top candidate for hard coatings. On the contrary, h-BN offers a significant anisotropy in its mechanical properties and therefore might locally help to produce micro-cracks parallel to the loading direction, resulting in toughening. (ii) c-BN and h-BN have different densities, with c-BN exhibiting the higher one. I.e., upon compression, a certain propensity of forming c-BN exists. However, this phase transformation seems not to be trivial, and it remains unclear if the c-BN to h-BN phase transformation could be exploited from a materials engineering viewpoint. While first indications of improved mechanical properties by using a c-BN / h-BN composite do exist [Kupczyk2023], the role of phase transformations to the best of our knowledge remains unknown for this system and could be an interesting topic for future evaluation.
@@
A number of material properties were discovered in perovskite photovoltaics that have not been considered with respect to photovoltaics in the past, but which may benefit the generation and extraction of photogenerated charge carriers and hence the power conversion efficiency of the devices. Among these properties, a ferroelectric in-plane polarization of methylammonium lead iodide (MAPbI3) was vividly discussed to help with the separation of electrons and holes as well as their conduction to the respective electrodes along ferroelectric domain walls [rossi_ferroelectric_2018, rohm_ferroelectric_2017]. So far, the ferroelectric properties are genuine to MAPbI3 solar cells and some derivatives thereof [schulz_ferroelectricity_2022]. The necessary precondition is a non-centrosymmetric unit cell of the crystal, which is why this concept has not played a role in any of the established solar technologies, such as silicon, so far [rohm_ferroelectric_2019, breternitz_role_2020]. Yet, it may well play a role in future novel perovskite-inspired materials [krishnaiah_perovskite-inspired_2025].
The requirement of crystallinity and non-centrosymmetric unit cells seemingly does not link to organic solar cells, which are mostly amorphous. Yet, the great design flexibility of organic molecules can be used to introduce asymmetric polar functional groups, which are already in use today in piezoelectric polymers for organic touch sensors, e.g., polyvinylidene fluoride (PVDF). If, in addition, ways are found to orient these molecules, an organic analogy to in-plane polarization may be designed, which then also benefits the charge carrier separation in organic solar cells. The challenge here is that in-plane polarization is notoriously difficult to control via external electrical poling since the electrodes in organic solar cells only allow for the application of an out-of-plane electric field. [roehm2020] Hence, in-plane polarization probably must be part of the conformation of the film itself through post-processing or by molecular self-assembly. Covalent organic frameworks (COFs) may also mimic a crystal structure, which then can be modified for in-plane polarization.
@@
[…]
[heading=bibintoc]

### v2

@@ Abstract
S6 Detailed discussion of interesting concept combinations
Below is a more detailed discussion of the 5 selected examples of new combinations of concepts shown in the results section of the main text.
@@
During the concept prediction phase of the AI model trained on open-access scientific abstracts, an interesting connection between two research areas emerged. The AI identified “conventional ceramics” and “graphene oxide” as a potentially promising combination. While the integration of graphene oxide with ceramics has been explored in numerous studies, these approaches typically rely on different synthesis methods and subsequent mixing. In contrast, the method presented here is both unique and simple, yielding a distinctive material: a nanometer-thin multilayer graphene scaffold coated with an approximately 200 thick iron oxide layer (see Extended Data Fig. 2). This AI-predicted research direction aligned with prior experimental work that had been conducted but never published, highlighting the model’s potential to identify overlooked but valuable material synthesis strategies.
@@
The resulting composite featured a highly conductive multilayer graphene scaffold coated with an approximately $200\text{\,}\mathrm{nm}$ thick layer of iron oxide. This material exhibited high electrochemical capacity and reversibility, making it a strong candidate for conversion-type electrode materials. The improved redox kinetics were attributed to the high electronic conductivity of graphene and the intimate contact with the thin iron oxide layer, enhancing the overall electrochemical performance.
@@
Recent research underscores that molecular tensile strain management might be regarded as a more general proxy for interface design. Brabec and Friederich reported in Science 2024 torsional flexibility of HTL molecules, which was represented by the introduction of TPA units, as one of several design criteria for efficient perovskite solar cells [31]. We note that the suggestion "tensile strain + molecular architecture" was generated in December 2023. Properties such as film formation and homogenous film quality benefited from that as well. In another currently ongoing study by Brabec et al. (in submission), we found that embedding plasticizing molecules dramatically increases the mechanical and operational stability of thin film solar cells.
@@
The link between stress-induced phase transformation and hexagonal boron nitride (h-BN) is best explained in the context of enhancing the fracture toughness of ceramics by phase transformations. For instance, in zirconia, a tetragonal-to-monoclinic phase transformation results in a volume expansion in the vicinity of a crack, which locally results in compressive stress both, deflecting and stopping crack growth. This concept is applied to other metastable phases in hard coatings [49]. A second concept of reaching an enormous tensile strength with simultaneously high fracture toughness, which both typically exclude each other, is using local anisotropy. This was shown in pearlite wires, where multiple micro-cracks parallel to the loading direction were formed and enhanced the fracture toughness substantially [50]. There are at least two arguments why thinking of stress-induced phase transformations and h-BN is reasonable. (i) Cubic boron nitride (c-BN), having a Wurtzite crystal structure, is used as a hard coating because of its enormous hardness. In contrast, the hardness of h-BN is significantly lower, rendering it not a top candidate for hard coatings. On the contrary, h-BN offers a significant anisotropy in its mechanical properties and therefore might locally help to produce micro-cracks parallel to the loading direction, resulting in toughening. (ii) c-BN and h-BN have different densities, with c-BN exhibiting the higher one. I.e., upon compression, a certain propensity of forming c-BN exists. However, this phase transformation seems not to be trivial, and it remains unclear if the c-BN to h-BN phase transformation could be exploited from a materials engineering viewpoint. While first indications of improved mechanical properties by using a c-BN / h-BN composite do exist [51], the role of phase transformations to the best of our knowledge remains unknown for this system and could be an interesting topic for future evaluation.
@@
A number of material properties were discovered in perovskite photovoltaics that have not been considered with respect to photovoltaics in the past, but which may benefit the generation and extraction of photogenerated charge carriers and hence the power conversion efficiency of the devices. Among these properties, a ferroelectric in-plane polarization of methylammonium lead iodide (MAPbI3) was vividly discussed to help with the separation of electrons and holes as well as their conduction to the respective electrodes along ferroelectric domain walls [32, 33]. So far, the ferroelectric properties are genuine to MAPbI3 solar cells and some derivatives thereof [35]. The necessary precondition is a non-centrosymmetric unit cell of the crystal, which is why this concept has not played a role in any of the established solar technologies, such as silicon, so far [35, 36]. Yet, it may well play a role in future novel perovskite-inspired materials [51].
The requirement of crystallinity and non-centrosymmetric unit cells seemingly does not link to organic solar cells, which are mostly amorphous. Yet, the great design flexibility of organic molecules can be used to introduce asymmetric polar functional groups, which are already in use today in piezoelectric polymers for organic touch sensors, e.g., polyvinylidene fluoride (PVDF). If, in addition, ways are found to orient these molecules, an organic analogy to in-plane polarization may be designed, which then also benefits the charge carrier separation in organic solar cells. The challenge here is that in-plane polarization is notoriously difficult to control via external electrical poling since the electrodes in organic solar cells only allow for the application of an out-of-plane electric field. [37] Hence, in-plane polarization probably must be part of the conformation of the film itself through post-processing or by molecular self-assembly. Covalent organic frameworks (COFs) may also mimic a crystal structure, which then can be modified for in-plane polarization.
@@
S7 Dataset
Filtering
Cleaning
Supplementary Fig. 21: Overview of the dataset generation: (a) fetch publications from OpenAlex venues related to materials sciences, (b) filtering based on title and abstract (c) removal of selected characters.
Chemical Formulas
  1. 1.
    The subsequent word is another element
  2. 2.
    The subsequent word is an integer together with a + or -
  3. 3.
    The subsequent word is an integer, a + or a -
  4. 4.
    The subsequent word is a decimal
  5. 5.
    The subsequent word is a composition of elements of the periodic table
S8 LLM Fine-tuning
Hugging Face’s Python library was used to fine-tune the LLaMa-2-13B model [41, 52]. We prepare our annotated data by constructing a single string per data point as follows <s> + abstract + "#KEYWORDS#" + concepts + </s> where concepts corresponds to the string version of a Python list, e.g. [’brittle epoxy’, ’butadiene nitrile liquid rubber’, ...]. <s> and </s> represent the beginning of sequence (BOS) and, respectively, end of sequence (EOS) token for the LLaMa models. #KEYWORDS# serves as an own custom delimiter to signal the model when concept extraction should start. For the LLaMa tokenizer, we set pad_token_id = 0, bos_token_id = 1, and eos_token_id = 2. Being a CausalLM, we set the labels equal to the input token ids. Our LoraConfig is r = 16 (Lora attention dimension), lora_alpha = 32, target_modules=["q_proj", "v_proj"] (modules where the adapter is applied), and lora_dropout = 0.01 (dropout probability) [44]. Training took around an hour on a single A100 GPU.
S9 Hyperparameters of the link prediction models
Supplementary Table 9: Optimized hyperparameters of link prediction models.[a]
Model Layers Neurons per layer Learning rate
Baseline 7 [20, 300, 180, 108, 64, 10, 1] $5\cdot 10^{-4}$
Concept embeddings 5 [1536, 1024, 819, 10, 1] $1\cdot 10^{-3}$
Combination of features 7 [1556, 1556, 933, 559, 335, 10, 1] $1\cdot 10^{-3}$
The GNN link prediction model uses a two-layer GraphSAGE encoder with a hidden dimension of 256 and produces 128-dimensional node embeddings. These embeddings are then fed into an MLP decoder that also has a hidden dimension of 256 and operates on 128-dimensional inputs. Neighborhood sampling is performed with fanout $(k_{1},k_{2})=(20,15)$ for the first and second hop, respectively. The model is trained with a learning rate of $1\cdot 10^{-5}$; additionally, a ReLU activation function, dropout with $p_{\text{dropout}}=0.1$ is used.